OpenStudy (anonymous):

You had $130 to go shopping. You bought 2 pairs of shoes that costs $25 for 1 pair, 1 pair of pants at $29.99 and 3 shirts, each at the same price. The tax was $10.44 and you received $3.57 for your change. How much does it cost for 1 shirt?

OpenStudy (anonymous):

let x is the price of a pairs of shoes=$25 y is the price of the shirt so we have: 130- (2x+29.99+3y+10.44)=3.57 130- [2(25)+29.99+3y+10.44)=3.57 130-(50+29.99+3y+10.44)=3.57 130-(3y+90.43)=3.57 130-3y-90.43=3.57 39.57-3y=3.57 39.57-3.57=3y 36=3y y=36/3 y=12 the price of the shirt is $12

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

you are welcome is that correct?

OpenStudy (anonymous):

idk yet lol